OpenStudy (anonymous):

square root of 176. simplify it please.

OpenStudy (rogue):

\[176 = 16 * 11\]\[\sqrt {176} = \sqrt {16 * 11} = \sqrt {16} * \sqrt {11} = 4\sqrt{11}\]

OpenStudy (anonymous):

ah yes thank you all i need was the 16 x 11 part thank you very much

OpenStudy (precal):

look for perfect squares

OpenStudy (precal):